8 men can build a house in 60 hours. how many men does it take to build the house in 32 hours?
1
Expert's answer
2012-06-05T09:26:27-0400
let x men build the house in 32 hours then 8/x=32/60 so x=8*60/32=15
so 15
men build the house in 32 hours